# Adaptive Human Trajectory Prediction via Latent Corridors

### v1

@@ 5 Experimental Setup
Figure 3: Qualitative results on MotSynth (top; synthetic) and MOT and WildTrack (bottom; real). These examples show scenarios where our LC + per-scene finetune ATP method (orange) outperforms the scene-aware baseline (purple). In several MOTSynth examples, the baseline predicts the pedestrian floats into the air (top row), while our method has gained awareness of where the 3D ground plane lies in the 2D image. We also note that patterns of behaviour such as walking on the sidewalk instead of into the road (second row left) and walking up the traversable portion of stairs (second row right) is captured. On real data, we observe similar awareness of the ground plane and obstacles, as well as a better understanding of nuanced human behavior patterns such as crossing diagonally across a plaza.
Figure 4: Qualitative results over time. Our method’s predictions that saw a short number of human seconds in training ($2\%$) are shown in light orange, with a human seconds training time of up to $80\%$. With the latent corridor trained on a tiny amount of data, the predictions sometimes significantly improve, but other times are close to the baseline. When more human seconds of data have been seen, the adaptation results consistently improve.
Figure 5: Learning latent corridors over time. (a and b): The x-axis represents adaptation time in person-seconds, or the amount trajectories use to train $\mathcal{A}(\mathcal{P})$, and the y-axis represents the ADE. Results are normalized per-scene and averaged over models trained on 25 MOTSynth scenes (a) and 7 from MOT and WildTrack (b), with shaded area $\sigma/10$. On all data, even with a short adaptation time, our methods improve on the baselines, and as adaptation time increases, performance improves. Latent corridors + per-scene finetuning has the best performance. c) Comparison to baseline over many MOTSynth scenes for models trained with $8\%$ (point) and $80\%$ (arrowhead) h-s datasets. Each arrow represents one scene, with the ADE using our ATP method plotted against the scene-aware baseline ADE. For some deployment scenes, the scene-aware baseline suffices, whereas other scenes see much more significant benefits from our method.
We study our approach on synthetic and real datasets. Here, we detail these datasets, describe how we evaluate adaptation quality over time, and outline our trajectory predictor baselines.
@@ 5.1 Datasets
MOTSynth. We start in simulation with MOTSynth [12], a synthetic pedestrian detection and tracking dataset of over $700$ scenes with varying camera viewpoints and outdoor environments, with each video $90$ seconds long. Pedestrians carry out simple actions such as walking, standing, or running, and follow manually pre-planned flows as well as a collision avoidance algorithm. MOTSynth has over $17$ hours of video; we select a subset of approximately $11$ hours of video corresponding to the $437$ scenes with a static camera. Due to the large size of the dataset and perfect ground truth detections, MOTSynth was a good starting point.
@@
EarthCam. We evaluate our approach in-the-wild on data from https://www.earthcam.com/, which has livestream webcam data from around the world available, and from which we scrape 5 minute segments of data.
@@ 5.2 Train-Test Split Over Time in Human Seconds
Given a video of pedestrians in motion, we prepare our dataset to adapt to the scene as follows. We use provided annotations or run ByteTrack [48] to get tracklets of all $N$ detectable identities in the video. Tracklets are downsampled to 2 timesteps per second, then windowed by $20$ timesteps with an observed length $H=8$ and future length $T=12$. Then, we sort the $N$ identities chronologically by their first appearance in the scene. We select the first $80\%$ of identities that appear in the scene for our training dataset, and hold out the last $20\%$ for the testing set. When we conduct experiments over time, the testing set is always the same $20\%$ of identities, whereas the training set consists of the first $m\%$ of identities, $m\leq 80\%$. We report adaptation time in person-seconds of observation, where 1 person for 1 second is 1 person-second, so for instance, 30 people observed for 10 seconds each would be 300 person-seconds.
@@ 5.3 Base Predictor Pre-Training Dataset
We first pre-train our base predictor $\mathcal{P}$ described in Sec 4.1 on dataset $\mathcal{D}$ which consists of simulated human pedestrian agents moving around various outdoor environments, captured from a single static RGB camera with any point of view (not necessarily birds eye view). This dataset comes from 90-second MOTSynth videos with a static camera. We pre-train our predictor on a train-test split over time of $\mathcal{D}$ as described in the previous section.
@@ 5.4 Baselines
    Learned trajectory: Simplified PECNet [27], which is a learned predictor with position history but no scene.
@@
    Scene-aware: Simplified YNet [28], which is a learned predictor with position and scene input.
@@
We implement the learned trajectory baseline by taking the encoder, decoder and trajectory loss from PECNet [27]. This removes multimodality for simplicity, but uses the an architecture proven to be successful on the trajectory prediction problem. Similarly, YNet’s [28] encoder and trajectory decoder are used for the scene-aware baseline.
@@ 5.5 Metrics
Our experiments are evaluated using the Average Displacement Error (ADE) [32] and Final Displacement Error (FDE) [2] metrics. ADE is the average $l_{2}$ error between the entire predicted and ground truth trajectory, while FDE is the $l_{2}$ error between just the final point of the predicted trajectory and the final point of the ground truth trajectory.
@@ 6 Results
Method MOTSynth MOT Wildtrack NoLA
@@
Constant velocity 78.5 160.4 47.7 99.3 44.9 90.1 42.4 82.8
Learned trajectory (PECNet-Ours) 51.2 100.0 49.7 103.4 43.8 83.1 36.9 65.9
Scene-aware (YNet-Ours) 47.3 96.5 44.2 91.0 33.6 67.7 34.0 65.1
ATP (LC) 44.6 90.2 43.0 89.2 32.9 65.9 33.5 63.1
@@
ATP (LC + Per-Scene Finetune) - - 37.4 74.3 27.4 54.7 26.1 47.4
Table 1: Our method vs baselines on real-world datasets. The average ADE and FDE are in pixel space of (left-to-right) 437 MOTSynth scenes, 4 MOT scenes, 3 WildTrack scenes, and 2 NoLA webcam scenes. Across these synthetic and real-world deployment scenes, our latent corridor adaptation method is comparable to finetuning the predictor head and outperforms non-adaptive baselines. Our method with latent corridors and per-scene finetuning consistently outperforms all other approaches.
@@ 6.1 MOTSynth Results
We first evaluate if latent corridors enable predictor improvement when the deployment scene is in the pre-training dataset, $\mathcal{D}$. We train the baselines and two variants of our ATP method (LC and LC + Joint Finetune, described in Sec. 4.3) on the MOTSynth pre-training dataset $\mathcal{D}$. All models see the same training data, and both the scene-aware baseline and our adaptive method are conditioned on each scene semantic segmentation map. Results are in the first column of Table 1. The learned trajectory baseline significantly improves over constant velocity, and adding in scene awareness results in a $7.6\%$ performance gain on ADE and $3.5\%$ gain on FDE. Our latent corridor approach results in a $5.7\%$ and $9.9\%$ improvement over the scene-aware baseline for ADE, without and with joint finetuning of the predictor layer respectively, and a $6.5\%$ and $11.6\%$ improvement on FDE. This indicates that our latent corridor approach is able to more effectively learn scene-conditioned information that is useful for the trajectory prediction task.
@@
Next, for a random subset of $25$ of the MOTSynth scenes, we additionally train latent corridors with per-scene finetuning (LC + Per-Scene FT, from Sec 4.3). The datasets $\mathcal{D}^{\prime}_{0:t}$ correspond to increasing human-second lengths. The test set trajectories are the last $20\%$ of agents in the deployment scene, and the training sets consist of the first $2,4,8,16,32,48,64$ and $80\%$ of agents to enter the scene. Results normalized per-scene and averaged are shown in Fig. 5b. Using only latent corridors has comparable results to only finetuning the last layer for each deployment scene, while latent corridors with per-scene finetuning yields significant performance gains. We visualize the ADE results per-scene with our LC + finetuning method trained with $8\%$ and then trained with $80\%$ of the data in Fig. 5c. We see that while the effectiveness of the adaptation for varying time horizons differs for each deployment scene, there are many scenes where adaptation yields significant gains, and some where the improvement on ADE error is up to $63.4\%$. We hypothesize that the scenes where our method does not help exhibit behaviors and environment geometries that the prior $\mathcal{P}$ is sufficient for.
@@
Qualitative results. Visualizations from the per-scene models can be seen in Fig. 3 and Fig. 4. We observe many examples where the scene-aware baseline seems to lack awareness of the ground plane. There are multiple instances where the scene-aware baseline predicts a pedestrian floats into the air (Fig. 3 top row and Fig. 4), whereas our predictions lie closely in line with the ground truth trajectory in terms of distance from the ground plane. This holds even for non-planar ground planes: When a pedestrian walks up stairs, our model seems to better understand their 3D structure, for example in the bottom right of Fig. 4. Additionally, our approach captures trends of behaviour in scenes such as walking on the sidewalk instead of into the road, walking around the rotunda, and walking up the section of smaller, easily traversable steps (see Fig. 3 second row, left to right).
@@ 6.2 MOT and WildTrack Results
We next investigate if latent corridors help predictor adaptation when the deployment scene is outside of the pre-training dataset, and if latent corridors outperform ATP via direct finetuning on deployment data. Specifically, we use the base predictor $\mathcal{P}$ which only saw MOTSynth data and then directly do sim-to-real adaptation via latent prompting, finetuning, or a combination thereof on seven real human pedestrian scenes from MOT and Wildtrack. Plots of the ADE over time averaged over these scenes is shown in Figure 5a. Regardless of the ATP setting, adaptation over time results in a consistent error reduction compared to the baselines. While only finetuning seems to be slightly more effective on real-world scenes than prompting alone, our latent corridors + per-scene finetuning approach is significantly more effective ($11.2\%$) than finetuning alone. Quantitative results for these experiments can be seen in Table 1.
@@
6.3 EarthCam NoLA Results
Figure 6: Results on webcam data of New Orleans. We train our method on one intersection during the daytime, when pedestrians must obey vehicles moving through the street, and at nighttime, when pedestrians take over. During the day, the model learns to account for humans walking through crosswalks instead of diagonally into the intersection, and to account for transient push-carts in the area. We also see awareness of the 3D ground plane.
Finally, we evaluate the performance of our approach on truly “in-the-wild” data by scraping two 5-minute videos of pedestrian data captured from Bourbon Street in New Orleans: one video during daytime activity (top row, Fig. 6) and one video during nighttime activity (bottom row, Fig. 6). Additional results on different scenes of webcam data are in Sec 10. Quantitatively, the results on this in-the-wild data align with our results on real and synthetic data: our ATP model with latent corridors and per-scene finetuning outperforms the non-adaptive baselines and pure finetuning. Qualitatively, our method adapts to daytime patterns where pedestrians navigate around carts and don’t frequently walk diagonally through streets because of through-traffic. When the adaptive predictor observes nighttime behavior, it no longer has to respect these daytime patterns and learns to predict pedestrians as crossing the diagonally. Similarly to the prior datasets we explored, our ATP model again learns to ground pedestrian future behavior in the 3D ground plane.
@@ 7 Conclusion
In this work, we formalize and study the problem of adaptive trajectory prediction: the ability of human predictors to adapt to changing deployment conditions and environments. To this end, we proposed a lightweight adaptation approach grounded in image-based prompt tuning called latent corridors. Through extensive experiments on both simulated and real world pedestrian datasets, we observed that latent corridors enable a data-efficient way to adapt pre-trained predictors to new scene-specific human behavior.
Supplementary Material
In the supplementary material, we provide an ablation on the implementation of our prompting method and a visual overview of our ATP problem formulation. We also show additional qualitative results on MOTSynth and EarthCam data, as well as more detailed quantitative results.
@@ 8.1 Adaptive Trajectory Prediction Visualization
We provide an illustrative overview of adaptive trajectory prediction problem, formulated in Sec. 3, in Fig. 7.

### v2

@@ 5 Experimental Setup
We study our approach on synthetic and real datasets. Here, we detail these datasets, describe how we evaluate adaptation quality over time, and outline our trajectory predictor baselines. Together, our experiments on MOT, WildTrack and EarthCam scenes cover a diverse range of key real-world properties including different lighting conditions, flat vs varied environment topologies, different crowd densities, and a variety of types of scenes.
@@ 5.1 Datasets
MOTSynth. We start in simulation with MOTSynth [12], a synthetic pedestrian detection and tracking dataset of over $700$ $90$-second videos with varying camera viewpoints and outdoor environments. Pedestrians carry out simple actions such as walking, standing, or running, and follow manually pre-planned flows as well as a collision avoidance algorithm. MOTSynth has over $17$ hours of video; we select a subset of approximately $11$ hours of video corresponding to the $437$ scenes with a static camera. Due to the large size of the dataset and perfect ground truth detections, MOTSynth was a good starting point.
@@
EarthCam. We evaluate our approach in-the-wild on data from https://www.earthcam.com, which has livestream webcam data from around the world available, and from which we scrape four 5 minute segments of data.
@@ 5.2 Train-Test Split Over Time in Human Seconds
Given a video of pedestrians in motion, we prepare our dataset to adapt to the scene as follows. We use provided annotations or run ByteTrack [55] to get tracklets of all $N$ detectable identities in the video. Tracklets are downsampled to 2 timesteps per second, then windowed by $20$ timesteps with an observed length $H=8$ and future length $T=12$. Then, we sort the $N$ identities chronologically by their first appearance in the scene. We select the first $80\%$ of identities that appear in the scene for our training dataset, and hold out the last $20\%$ for the testing set. When we conduct experiments over time, the testing set is always the same $20\%$ of identities, whereas the training set consists of the first $m\%$ of identities, $m\leq 80\%$. We report adaptation time in human-seconds of observation, where 1 person for 1 second is 1 human-second, so for instance, 30 people observed for 10 seconds each would be 300 human-seconds.
@@ 5.3 Base Predictor Pre-Training Dataset
We first pre-train our base predictor $\mathcal{P}$ described in Sec 4.1 on a dataset $\mathcal{D}$ which consists of simulated human pedestrian agents moving around various outdoor environments, captured from a single static RGB camera with any point of view (not necessarily birds eye view). This dataset comes from 437 90-second MOTSynth videos with a static camera. We pre-train our predictor on a train-test split over time of $\mathcal{D}$ as described in the previous section.
@@ 5.4 Baselines
    Learned trajectory: Simplified PECNet [28], which is a predictor learned from position histories but no scene information.
@@
    Scene-aware: Simplified YNet [27], which is a learned predictor with position and scene input. See Supplement Sec. 3 for a comparison with YNet.
@@
We implement the learned trajectory baseline by taking the encoder, decoder and trajectory loss from PECNet [28]; we use a successful architecture on the trajectory prediction problem, removing multimodality for simplicity. Similarly, YNet’s [27] encoder and trajectory decoder are used for the scene-aware baseline.
@@ 5.5 Metrics
Our experiments are evaluated using the Average Displacement Error (ADE) [35] and Final Displacement Error (FDE) [2] metrics. ADE is the average $l_{2}$ error between the entire predicted and ground truth trajectory, while FDE is the $l_{2}$ error between just the final point of the predicted and ground truth trajectories.
@@ 6 Results
Figure 3: Qualitative results on MotSynth (top; synthetic) and MOT and WildTrack (bottom; real). These examples show scenarios where our LC + per-scene finetune ATP method (orange) outperforms the scene-aware baseline (purple). In several MOTSynth examples, the baseline predicts the pedestrian floats into the air (top row), while our method has gained awareness of where the 3D ground plane lies in the 2D image. We also note that patterns of behaviour such as walking on the sidewalk instead of into the road (second row left) and walking up the traversable portion of stairs (second row right) are captured. On real data, we observe similar awareness of the ground plane and obstacles, as well as a better understanding of nuanced human behavior patterns such as crossing diagonally across a plaza.
Figure 4: Qualitative results over time. Our method’s predictions trained on a short number of human seconds ($2\%$) are shown in light orange, to dark orange for a human seconds training time of $80\%$. With the latent corridor trained on a tiny amount of data, the predictions can significantly improve, but at times are close to the baseline. When more human seconds of data are seen, the adaptation results consistently improve.
Figure 5: Learning latent corridors over time. (a and b): The x-axis represents adaptation time in human-seconds, or the amount trajectories used to train $\mathcal{A}(\mathcal{P})$, and the y-axis represents the ADE. Results are normalized per-scene and averaged over models trained on 25 MOTSynth scenes (a) and 7 from MOT and WildTrack (b), with shaded area $\sigma/10$. On all data, even with a short adaptation time, our methods improve on the baselines, and as adaptation time increases, performance improves. Latent corridors + per-scene finetuning has the best performance. c) Comparison to baseline over many MOTSynth scenes for models trained with $8\%$ (point) and $80\%$ (arrowhead) human-second datasets. Each arrow represents one scene, with the ADE using our ATP method plotted against the scene-aware baseline ADE. For some deployment scenes, the scene-aware baseline suffices, whereas other scenes see much more significant benefits from our method.
Method MOTSynth MOT Wildtrack EarthCam
@@
Constant velocity 78.5 160.4 47.7 99.3 44.9 90.1 27.3 52.2
Learned trajectory (PECNet-Ours) 51.2 100.0 49.7 103.4 43.8 83.1 34.3 55.5
Scene-aware (YNet-Ours) 47.3 96.5 44.2 91.0 33.6 67.7 23.9 44.4
ATP (Finetune) - - 41.8 86.9 31.7 63.5 21.9 38.7
ATP (LC) 44.6 90.2 43.0 89.2 32.9 65.9 23.4 42.8
@@
ATP (LC + Per-Scene Finetune) - - 37.4 74.3 27.4 54.7 17.5 30.8
Table 1: Our method vs baselines on synthetic and real-world datasets. The average ADE and FDE in pixel space of (left-to-right) 437 MOTSynth scenes, 4 MOT scenes, 3 WildTrack scenes, and 4 EarthCam scenes. Across these synthetic and real deployment scenes, our latent corridor-only adaptation method is comparable to finetuning the predictor head and outperforms non-adaptive baselines. Our method with latent corridors and per-scene finetuning consistently outperforms all other approaches.
@@ 6.1 MOTSynth Results
We first evaluate predictor improvement enabled by latent corridors when the deployment scene is in the pre-training dataset, $\mathcal{D}$. We train the baselines and two variants of our ATP method (LC and LC + Joint Finetune, described in Sec. 4.3) on the MOTSynth pre-training dataset $\mathcal{D}$. All models see the same training data, and both the scene-aware baseline and our adaptive method are conditioned on scene semantic segmentation maps. Results are in the first column of Table 1. The learned trajectory baseline significantly improves over constant velocity, and adding in scene awareness results in a $7.6\%$ performance gain on ADE and $3.5\%$ gain on FDE. Our latent corridor approach results in a $5.7\%$ and $9.9\%$ improvement over the scene-aware baseline for ADE, without and with joint finetuning of the predictor layer respectively, and a $6.5\%$ and $11.6\%$ improvement on FDE. This indicates that our latent corridor approach more effectively learns scene-conditioned information that is useful for the trajectory prediction task.
@@
Next, for a random subset of $25$ of the MOTSynth scenes, we additionally train latent corridors with per-scene finetuning (LC + Per-Scene FT, Sec 4.3). The datasets $\mathcal{D}^{\prime}_{0:t}$ correspond to increasing human-second lengths. The test set trajectories are the last $20\%$ of agents in the deployment scene, and the training sets consist of the first $2,4,8,16,32,48,64$ and $80\%$ of agents to enter the scene. Results normalized per-scene and averaged are shown in Fig. 5b. Using only latent corridors performs comparably to only finetuning the last layer for each deployment scene, while latent corridors with per-scene finetuning yields significant performance gains. We visualize the ADE results per-scene with our LC + finetuning method trained with $8\%$ and then trained with $80\%$ of the data in Fig. 5c. We see that while the effectiveness of the adaptation for varying time horizons differs for each deployment scene, there are many scenes where adaptation yields significant gains, and some where the improvement on ADE error is up to $63.4\%$. We hypothesize that scenes where our method yields smaller gains exhibit behaviors and environment geometries that the prior $\mathcal{P}$ is sufficient for.
@@
Qualitative results. Visualizations from the per-scene models can be seen in Fig. 3 and Fig. 4. We observe many examples where the scene-aware baseline seems to lack awareness of the ground plane. In multiple instances, the scene-aware baseline predicts a pedestrian floats into the air (Fig. 3 top row and Fig. 4), whereas our predictions lie closely in line with the ground truth trajectory in terms of distance from the ground plane. This holds even for non-planar ground planes: When a pedestrian walks up stairs, our model seems to better understand their 3D structure, for example in the bottom right of Fig. 4. Additionally, our approach captures trends of behaviour in scenes such as walking on the sidewalk instead of into the road, walking around the rotunda, and walking up the section of smaller, easily traversable steps (see Fig. 3 second row, left to right).
@@ 6.2 MOT and WildTrack Results
We next investigate if latent corridors help predictor adaptation when the deployment scene is outside of the pre-training dataset, and if latent corridors outperform ATP via direct finetuning on deployment data. Specifically, we use the base predictor $\mathcal{P}$ which only saw MOTSynth data $\mathcal{D}$ and then directly do sim-to-real adaptation via latent prompting, finetuning, or a combination thereof using each of seven real human pedestrian scenes from MOT and Wildtrack as $\mathcal{D}^{\prime}$. Plots of the ADE over time averaged over these scenes is shown in Figure 5a. Regardless of the ATP setting, adaptation over time results in a consistent error reduction compared to the baselines. While only finetuning seems to be slightly more effective on real-world scenes than prompting alone, our latent corridors + per-scene finetuning approach is significantly more effective ($11.2\%$) than finetuning alone (see Table 1).
@@
6.3 EarthCam Results
Figure 6: Results on EarthCam webcam data. At a cafe in Jamaica (top left 2 panels) with an overlook and stairs on the edge of the water, our model (orange) is able to correctly predict that people will stick to moving up and down the path created by the stairs, while the scene-aware baseline (purple) predicts people will walk directly into the water or over the edge of the path. In Times Square, a billboard depicts human actors in motion (top right 2 panels). The scene-aware baseline assumes that these actors will move following their observed history, while our model correctly predicts that anyone in the area of the billboard will stay in the billboard. We train our method on one NoLA intersection during the daytime (bottom left 2 panels), when pedestrians must obey vehicles moving through the street, and at nighttime (bottom right 2 panels), when pedestrians take over. During the day, the model learns to account for humans walking through crosswalks instead of diagonally into the intersection, and to account for transient push-carts in the area.
Finally, we evaluate the performance of our approach on truly “in-the-wild” data by scraping four 5-minute videos of pedestrian data captured in different locations: Rick’s Cafe in Jamaica, Times Square in New York City, and Bourbon Street in New Orleans during daytime and nighttime. We use ATP as in Sec. 6.2. Quantitatively, the results on this in-the-wild data align with our results on real and synthetic data: our ATP model with latent corridors and per-scene finetuning outperforms the non-adaptive baselines and pure finetuning (see Table 1). Qualitatively in Fig. 6, we see a variety of interesting adaptations. At the cafe, there is a complicated path through an overlook that twists down stairs towards the water. The scene-aware baseline often predicts that people will jump over a ledge, whereas our method learns the boundaries of the paths that people follow and is able to predict that people will stay within those boundaries. In Times Square, the scene-aware baseline does not recognize that people on a billboard will stay within the billboard, whereas our method is able to recognize that. In the NoLA videos, our method adapts to daytime patterns where pedestrians navigate around carts and don’t frequently walk diagonally through streets because of through-traffic, but when the adaptive predictor observes nighttime behavior, it no longer has to respect these daytime patterns and learns to predict pedestrians as crossing diagonally. Similarly to the prior datasets, our ATP model again learns to ground pedestrian future behavior in the 3D ground plane.
@@ 7 Conclusion
In this work, we formalize and study the problem of adaptive trajectory prediction: the ability of human predictors to adapt to changing deployment conditions and environments. To this end, we proposed a lightweight adaptation approach grounded in image-based prompt tuning called latent corridors. Through extensive experiments on both simulated and real-world pedestrian datasets, we observed that latent corridors enable a data-efficient way to adapt pre-trained predictors to new deployment-scene-specific human behavior.
Acknowledgements
We thank Anastasios Angelopoulos, Antonio Loquercio, and Jathushan Rajasegaran for useful discussions and feedback. This work was supported by ONR MURI N00014-21-1-2801 and a NSF Graduate Fellowship.
Supplementary Material: Adaptive Human Trajectory Prediction via Latent Corridors
Neerja Thakkar Karttikeya Mangalam Andrea Bajcsy Jitendra Malik
In the supplementary material, we provide an ablation on the implementation of our prompting method and a visual overview of our ATP problem formulation. We also show additional qualitative results on MOTSynth and more detailed quantitative results on EarthCam data, quantitative results on ETH/UCY, and compare our implementation of the scene-aware baseline to the original YNet paper [27]. Video results can be seen at this webpage.
@@ 8.1 Adaptive Trajectory Prediction Visualization
We provide an illustrative overview of adaptive trajectory prediction problem, formulated in main text Sec. 3, in Fig. 7.
8.2 ATP via Latent Corridors on Architectures Beyond YNet
ATP is an architecture-agnostic paradigm, and latent corridors are also not specific to Y-Net but rather can also work on different architectures. To demonstrate this, we experimented with the Learned Trajectory (PECNet-Ours) architecture on the 473 MOTSynth scenes. Taking the pretrained PECNet-Ours model, we summed a per-scene 16 parameter latent directly to the input, eight $xy$ coordinates. Training the 16D latent along with finetuning the last layer of PECNet-Ours, we see an improvement of $10.2\%$ on ADE and $10.4\%$ on FDE.
Method ADE FDE
Learned Trajectory (PECNet-Ours) 51.2 100.0
ATP (LC + Joint Finetune) 46.0 89.6
Table 2: Results of applying ATP via latent corridors to PECNet. ATP using 16D latent corridors with joint finetuning improves the performance of PECNet.